# Gaussian in the Wild: 3D Gaussian Splatting for Unconstrained Image Collections

### v1

@@ 5.4 Appearance tuning experiment
Since we explicitly model the scene’s appearance as unchanged intrinsic and varying dynamic features, we can adjust the impact of dynamic appearance features on intrinsic appearance by multiplying $df_{i}$ by a proportional weight. We also compare it with Ha-NeRF and CR-NeRF by applying the same weights to extracted image features. Qualitative results are shown in Fig. 5. Ha-NeRF and CR-NeRF show odd coloration in the sky and buildings at low weights, while at high weights, they fail to capture the detailed highlights and enhance the style of sufficient illumination, resulting in darker building colors. In contrast, our method aligns more closely with human understanding of the physical world. As the weight increases from small to large, our method gradually applies extracted environmental impacts to the scene, e.g. appearing highlights on pillars and enhancing illumination. This demonstrates that explicitly separating intrinsic and dynamic appearances aids the model in clearly learning and distinguishing between the two, thereby achieving more flexible tuning over the scene’s appearance.
@@ 5.5 Limitations
While GS-W outperforms previous methods, it still has limitations. It struggles with capturing complex lighting variations and reconstructing specular reflections. Additionally, it faces challenges in accurately reconstructing textures in frequently occluded scenes, such as the floor texture in the Brandenburg Gate scene. Future research may focus on developing new appearance modeling techniques to address these issues.
@@ A Video demo
We strongly recommend readers to watch the video demo provided in the supplementary materials. The video showcases the novel view synthesis achieved by GS-W across various scenes, based on different reference images. It also illustrates the variations in scene appearance attained through the interpolation of different dynamic features $df$. Additionally, we provide a visual comparison between GS-W and CR-NeRF in terms of novel view synthesis and appearance tuning. The video illustrates GS-W’s superior reconstruction of scene geometry and precise capture of environmental factors from the reference images, maintaining multi-view consistency. Furthermore, it highlights GS-W’s ability to adjust scene appearance in a manner more aligned with human perception by weighting the features extracted from the reference images.
@@ C Further experiments
C.1 K value selection
In our method statement, K represents the number of feature maps used for sampling. We conducted experiments under the Sacre Coeur scene to determine an effective K, setting K to 1, 2, 3, and 4 respectively. The experimental results are presented in Tab. 4. We observe that the performance on the test set does not improve when K exceeds 3 while bringing more computation cost. Thus, we reasonably set K to 3.
C.2 More appearance tuning results
@@
C.3 Synthetic Lego dataset
@@
Figure 8: Style transfer from unseen images or across buildings for the three scenes. The first row represents the reference images used for extracting appearance features, while the subsequent three rows depict images rendered based on the views of the content images and the corresponding appearance features of the reference images. This demonstrates that our method can effectively perform style transfer.
@@ D Style transfer
Similar to CR-NeRF, our method can also transfer style from images to the scene, as shown in Fig. 8. However, for style transfer across buildings or from unseen images, the absence of camera poses from the provided reference images makes it difficult to map Gaussian points to the projection feature map for sampling. Meanwhile, in style transfer, there lack of physical significance to using the projection sampling method. Therefore, when performing style transfer from unseen images, we set all features $f^{P}$ sampled from the projection feature map to 0. The results demonstrate that our method can effectively transfer the style provided by reference images to the scene.

### v2

@@ 5.4 Appearance tuning experiment
Since we explicitly model the scene’s appearance as unchanged intrinsic and varying dynamic features, we can adjust the impact of dynamic appearance features on intrinsic appearance by multiplying $df_{i}$ by a proportional weight. We also compare it with Ha-NeRF and CR-NeRF by applying the same weights to extracted image features. Qualitative results are shown in Fig. 5. Ha-NeRF and CR-NeRF show odd coloration in the sky and buildings at low weights, while at high weights, they fail to capture the detailed highlights and enhance the style of sufficient illumination, resulting in darker building colors. In contrast, our method aligns more closely with human understanding of the physical world. As the weight increases from small to large, our method gradually applies extracted environmental impacts to the scene, e.g. appearing highlights on pillars and enhancing illumination. This demonstrates the importance of dynamic features in capturing environmental information and explicitly separating intrinsic and dynamic appearances aid the model in clearly learning and distinguishing between the two, thereby achieving more flexible tuning over the scene’s appearance.
@@ 5.5 Limitations
While GS-W outperforms previous methods, it still has limitations. It struggles with complex lighting variations, specular reflections, and accurately reconstructing textures in frequently occluded scenes, like the floor texture in the Brandenburg Gate scene. Additionally, it assumes known image poses when incorporating appearance information from reference images. Future research may focus on developing new appearance modeling techniques to address these issues.
@@ A Video demo
We strongly recommend readers to watch the video demo provided in webpage. The video showcases the novel view synthesis achieved by GS-W across various scenes, based on different reference images. It also illustrates the variations in scene appearance attained through the interpolation of different dynamic features $df$. Additionally, we provide a visual comparison between GS-W and CR-NeRF in terms of novel view synthesis and appearance tuning. The video illustrates GS-W’s superior reconstruction of scene geometry and precise capture of environmental factors from the reference images, maintaining multi-view consistency. Furthermore, it highlights GS-W’s ability to adjust scene appearance in a manner more aligned with human perception by weighting the features extracted from the reference images.
@@ B.2 Initialization of sampling coordinates
Table 4: Average quantitative results on the Brandenburg, Sacre, and Trevi scenes.
[…]
global feature 22.43 0.8555 0.1349 56.4 123.0
K-planes-hybrid 22.40 0.7604 0.2731 0.67 133.3
Ours 24.70 0.8655 0.1242 50.7 122.7
Figure 6: More qualitative comparison results on Brandenburg. Our method captures local highlights and avoids artifacts caused by transient objects, as observed in K-Planes and 3DGS.
@@ C Further experiments
C.1 More comparison
Global appearance feature. To highlight the superiority of independently sampled dynamic appearance features over a single global feature, We replace each point’s dynamic appearance feature with a global one extracted from the reference image. As shown in Tab. 4 and Fig. 6, the global feature method achieves an inferior result and fails to recover local appearance details on the pillars and sky.
K-Planes. Aiming to further demonstrate the advantages of our method in rendering quality and speed, we compared it with K-Planes[10]. Unlike NeRF-based methods, K-Planes uses a hybrid 3D representation that achieves lower storage and faster rendering. Quantitative results, including metrics for storage and rendering efficiency, along with qualitative results, are presented in Tab. 4,Tab. 7,Fig. 6 and Fig. 9. Our method continues to achieve superior rendering quality and speed in comparison.
3DGS. In Fig. 6 and Fig. 9, we provide visual comparisons between 3DGS and our method. It’s clear that 3DGS fails to capture scene appearance variations and handle transient objects, leading to noticeable flaws and artifacts.
C.2 K value selection
In our method statement, K represents the number of feature maps used for sampling. We conducted experiments under the Sacre Coeur scene to determine an effective K, setting K to 1, 2, 3, 4, 8, and 16 respectively. The experimental results are presented in Tab. 5. We observe that the performance on the test set does not improve when K exceeds 3 while bringing more computation cost. Thus, we reasonably set K to 3.
C.3 More appearance tuning results
@@
Table 7: Quantitative results on the test set of four NeRF-OSR scenes. Our method outperforms other methods across all scenes on all metrics.
stjohann lwp st europa
[…]
3DGS 17.32 0.7430 0.313 15.43 0.7000 0.3360 15.23 0.5920 0.4240 16.32 0.6520 0.3270
K-planes-hybrid 20.39 0.7548 0.3366 21.65 0.7482 0.3397 19.66 0.6151 0.4361 18.75 0.6487 0.4549
NeRF-W 21.38 0.8200 0.1940 21.29 0.7610 0.2950 19.68 0.6310 0.4010 19.55 0.6870 0.3470
Ha-NeRF 19.93 0.7870 0.2100 21.32 0.7560 0.2780 20.56 0.6360 0.3780 18.76 0.6610 0.3480
CR-NeRF 22.27 0.8350 0.1750 22.61 0.7850 0.2580 21.67 0.6610 0.3600 19.92 0.6960 0.3100
GS-W (Ours) 26.23 0.8927 0.1133 24.44 0.8272 0.2082 22.45 0.6900 0.3147 22.04 0.7577 0.2309
Figure 9: Qualitative comparison results on the NeRF-OSR dataset test set are shown for the stjohann, lwp, st, and europa scenes. Our method reconstructs the scenes with greater detail compared to other methods.
Figure 10: Style transfer from unseen images or across scenes for the three scenes. The first row represents the reference images used for extracting appearance features, while the subsequent three rows depict images rendered based on the views of the content images and the corresponding appearance features of the reference images. This demonstrates that our method can effectively perform style transfer.
C.4 Synthetic Lego dataset
@@
C.5 NeRF-OSR dataset
We further evaluate the robustness of GS-W using the NeRF-OSR dataset[36]. We experiment with four scenes - europa, lwp, st, and stjohann - with 12.5% of images from each sequence as the test set. Each training set contains approximately 350 images, and each test set contains about 50 images.
We compared our method with NeRF-W, Ha-NeRF, CR-NeRF, 3DGS, and K-Planes, using their default settings. Evaluation results on the test set, shown in Tab. 7 and Fig. 9, demonstrate that our method consistently achieves the best performance. Other methods struggle to capture local details, and 3DGS fails to handle variations in scene appearance and transient occlusions. In contrast, our method accurately captures local details and effectively handles transient occlusions.
Despite our method achieving good performance, this dataset, not sourced from the internet, features limited appearance variations per scene. Consequently, the reconstructed scenes may exhibit inconsistencies in appearance across different viewpoints. In the future, training a more generalized dynamic appearance extractor across multiple scenes could effectively address and improve this issue.
@@ D Style transfer
Similar to CR-NeRF, our method can also transfer style from images to the scene, as shown in Fig. 10. However, for style transfer across scenes or from unseen images, the absence of camera poses from the provided reference images makes it difficult to map Gaussian points to the projection feature map for sampling. Meanwhile, in style transfer, there lack of physical significance to using the projection sampling method. Therefore, when performing style transfer from unseen images, we set all features $f^{P}$ sampled from the projection feature map to 0. The results demonstrate that our method can effectively transfer the style provided by reference images to the scene.